Grade 1210 grade science

Compute the heat generated while transferring 96000 coulomb of charge in one hour through a potential difference of 50 V.

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Structure


Charge = 96000 coulomb
Time = 1 hour = 3600 seconds
Potential difference = 50 V
I = Q/T
I = 96000/3600 = 80/3 amp
V = I × R
50 = 80/3 × R
Thus. R = 15/8 ohm
Heat = I2RT
= (80/3)2 × 15/8 × 3600
= 800 × 6000
= 4800000 joules.

Thanks